1.14 Consider events ArAg, Avon a sample space Ω. (a) Suppose A, c A-... c AN . Evaluate P(AIA)for i < j and for i > (b) Evaluate the set CAnd D1 A (c) Prove/Disprove: N-1 n AN ) = 1.
Answer #1

(a)

For i < j, P(Ai cap Aj) = P(Ai)

P(Ai | Aj) = P(Ai cap Aj) / P(Aj) = P(Ai) / P(Aj)

For i > j, P(Ai cap Aj) = P(Aj)

P(Ai | Aj) = P(Ai cap Aj) / P(Aj) = P(Aj) / P(Aj) = 1

(b)

C- A-A1 U A2 U.... U AnAn

D = igcap_{i=1}^{N}A_i = A_1 cap A_2 cap.... cap A_n = A_1

(c)

(i)

A_1 subset A_N implies A_1 and ar{ A_N} are disjoint sets.

Thus, Ai UANメΩ

(ii)

A_1 subset A_N implies A_1 and ar{ A_N} are disjoint sets.

Thus,

A_1 cap ar{ A_N} e phi

iii)

P(A_1 cup overline{overline{A_1} cap A_2}... cup overline{overline{A_{N-1}} cap A_N})

P(A1 U A1 U A2 U AN-1 U AN)

P(A1 U A3 U U AN-1 U A-... U AN)

PLAN-1 U A2) As, A_2 subset A_N Rightarrow overline{A_N} subset overline{A_2}

e 1